# The Polar Express: Optimal Matrix Sign Methods and Their Application to the Muon Algorithm

### v2

@@ 3   Approximations by Compositions of Polynomials
Iterative methods for $\operatorname*{polar}(\bm{M})$ can be seen in this light. For instance, the degree-5 Newton-Schulz method uses the polynomial update $p_{t}(x)=\frac{15}{8}x-\frac{10}{8}x^{3}+\frac{3}{8}x^{5}$ for each $t=1,\ldots,T$. The composition $p=p_{T}\circ\cdots\circ p_{1}$ approximates $\operatorname*{sign}(x)$, and the approximation error goes to 00 as $T$ grows. In this paper, we ask the following question: what choice of $p_{T}\circ\cdots\circ p_{1}$ gives the best approximation to $\operatorname*{sign}(x)$?
@@
In other words, the problem given in (6) reduces to that of finding a “uniform” or “minimax” approximation to the constant function $x\mapsto 1$ over the interval $[\ell,u]$, as given in (7). Uniform approximation on an interval by polynomials or rational functions of a given degree is a central topic in approximation theory; see e.g. [46]. Here, we seek an approximation of a particular form—a composition of odd polynomials of fixed degrees. In the next section, we solve the optimization problem of (7) and use the solution to create Polar Express. Figure 2 (a) shows the resulting $p^{*}$ polynomial labeled as PolarExp, as compared to the Jordan’s method in [22], and the six iterations of You’s method in [9].
@@ 4.1   Greedy is optimal
The key observation is that the polynomial used in each iteration can be chosen greedily, given the choice of polynomials from the previous iterations. For the first iteration, we choose $p_{1}$ so as to map the interval $[\ell,u]$ as close to $1$ as possible. That is, it minimizes $\max_{x\in[\ell,u]}|1-p_{1}(x)|$. The image of $p_{1}$ will be a new interval $[\ell_{2},u_{2}]$, where
@@
We now pick $p_{2}$ to map the interval $[\ell_{2},u_{2}]$ as close to $1$ as possible, obtaining a new interval $[\ell_{3},u_{3}]$ that is the image of $[\ell,u]$ through $p_{2}\circ p_{1}$. We continue this process for as many iterations as desired.
@@
Then the new error, lower and upper bounds can be computed through
@@
Furthermore, the composition $p^{\star}:=p_{T}\circ p_{T-1}\circ\cdots\circ p_{1}$ is optimal and the error is given by:
@@
Our theory translates entirely to this more general case. However, for simplicity we assume $d=d_{t}$ for all $t=1,\ldots,T$. Our setting is similar to that of [28], which considers the closely related problem of choosing the depth $T$ and degrees $d_{1},\ldots,d_{T}$ such that $p$ approximates $\operatorname*{sign}$ up to a prescribed error tolerance while minimizing the number of scalar multiplications. Interestingly, from [28, Table 2] the optimal choice of degrees is $d_{t}=5$ for almost all iterations. This justifies choosing $d$ to be a constant and our use of $d=5$ in particular.
Fortunately, (10) shows that once $p_{t}$ has been found, we can compute the new lower and upper bounds $\ell_{t+1}$ and $u_{t+1}$ and the approximation error simply by evaluating $p_{t}(\ell_{t})$. Hence, for any fixed upper and lower bounds on the singular values of $\bm{M}$, we can precompute the polynomials $p_{1},\ldots,p_{T}$ and the bounds $[\ell_{1},u_{1}],\ldots,[\ell_{T+1},u_{T+1}]$. Then, applying the iterative procedure of (5), the final iterate $\bm{X}_{T}$ will satisfy the following error bound
@@
In fact, Theorem 4.3 underestimates how fast our method converges. For degree $d=5$, our method converges about twice as fast as Newton-Schulz (compare with [11, Section 3.1]). Furthermore, the same analysis applies even if $p^{*}$ is constructed using a “lower bound” $\ell$ that was too high. That is, replacing $\ell$ on the right-hand side of (13) by $\sigma_{\min}$, the theorem holds even if $p^{*}$ is constructed to be optimal on the interval $[\ell,1]$ for $\ell>\sigma_{\min}$. Intuitively, when $\ell=u=1$, the polynomial $p^{*}$ coincides exactly with the Newton-Schulz method. Mistakenly setting $\ell>\sigma_{\min}$, we obtain a polynomial that converges slower than the optimal polynomial but faster than Newton-Schulz, so the guarantee of Theorem 4.3 still holds (cf. [11, Theorem 3.3]).
@@ 4.2   Finding the optimal polynomial for each iteration
Theorem 4.1 shows that we can solve (7) by greedily choosing the optimal approximation $p_{t}\in\mathbb{P}_{d}^{\operatorname*{odd}}$ for each interval $[\ell_{t},u_{t}]$ for $t=1,\ldots,T$. In this section, we show how to find each $p_{t}$. Since we are now focused on just one iteration, we drop the subscripts. Given $\ell$ and $u$, we wish to solve the following optimization problem:
@@
That is, we seek a minimax or uniform approximation of the function $x\mapsto 1$ on $[\ell,u]$ from the set of odd polynomials. (Equivalently, we seek a minimax optimal approximation to $\operatorname*{sign}(x)$ on $[-u,-\ell]\cup[\ell,u]$.)
Problems of the form (14) are well-studied in approximation theory and numerical analysis. The key mathematical insight underlying the solution is the Equioscillation Theorem, which we state formally for our setting in Lemma A.1. This theorem gives a surprising characterization of the optimal solution of (14): an odd $p$ is optimal for degree $12q+$ 1 if and only if there is a set of $q+2$ equioscillating points. This is a set of points at which $p$ achieves its maximum approximation error ±$\pm E$, and for which the sign of the error alternates. Even if the optimal approximation error $E$ is not known in advance, finding a set of $q+2$ equioscillating points for a given $E$ serves as a certificate that no better approximation error is achievable. The Equioscillation Theorem is the basis of the Remez algorithm [37, 38], a general tool that can be used to find (nearly) optimal polynomial approximations of a given degree to any function on any interval. With very minor modifications to handle the constraint that $p$ be odd, Remez can be used to directly solve (14).
@@
We briefly describe the solution for $d=3$. We seek a polynomial of the form $p(x)=ax+bx^{3}$. The Equioscillation Theorem stipulates that $p$ must have an equioscillating set of size 3. For $p$ to achieve its maximum error at a point $x$, $x$ must be a local extremum of $p(x)-1$ on the interval $[\ell,u]$. Thus, for $x$ to be eligible for membership in the equioscillating set, it must either be a true local extremum of $p(x)-1$ that happens to lie in $[\ell,u]$, or else one of the endpoints $\ell,u$. However, because $p$ is an odd cubic, it has at most one true local extremum on $\mathbb{R}_{\geq 0}$. Thus, to build an equioscillating set of three points, we must include $p$’s unique positive local extremum and both endpoints. This local extremum of $p$ occurs at $\sqrt{\frac{-a}{3b}}$. Therefore, we seek $a,b$ such that
@@
We now turn to the degree-5 case. The intuition of Algorithm 3 is as follows. For any fixed set of four points $\ell<q<r<u$, we can find an degree-5 odd polynomial $p(x)=ax+bx^{3}+cx^{5}$ that satisfies
@@
for some $E$ by solving a $4\times 4$ linear system in $a,b,c$ and $E$. Likewise, for any fixed degree-5 odd $p$, we can find its four (or fewer) local extrema on $[\ell,u]$ as follows: they occur at $\ell,u$ and the roots of $p^{\prime}$, which is an even degree-4 polynomial whose roots can easily be found by the quadratic formula. Algorithm 3 simply alternates between these two steps (solving for $a,b,c,E$ and solving for $q,r$) until the points $q,r$ converge. Once they have converged, $\{\ell,q,r,u\}$ forms an equioscillating set, so $p$ is the optimal polynomial. For more details, please see Appendix D.
@@ 4.4   Finite precision considerations
To fix this issue, we simply replace each polynomial $x\mapsto p_{t}(x)$ by $x\mapsto p_{t}(x/1.01)$. This safety factor corrects for round-off errors in previous iterations while only slightly changing the behavior of the polynomial on the interval $[\ell_{t},u_{t}]$, though it does cause the singular values to converge to $0.999998$ instead of to $1$. To correct for this, the safety factor can be omitted in the final iteration.
The second issue was identified in [34] and addressed in the context of polynomial iterations by Chen and Chow [11]. In general, iterative methods for $\operatorname*{polar}(\bm{M})$ aim to increase each singular value relative to the largest singular value; while $\sigma_{\min}(\bm{X}_{0})\ll\sigma_{\max}(\bm{X}_{0})$, after enough iterations, $\sigma_{\min}(\bm{X}_{t})\approx\sigma_{\max}(\bm{X}_{t})\approx 1$. However, the convergence of each singular value to $\sigma_{\max}$ may not be monotonic. Over the domain $[\ell_{t},u_{t}]$, our optimal polynomial $p_{t}$ oscillates repeatedly between $\ell_{t+1}$ and $u_{t+1}$, so some singular values that are near $u_{t}$ may get mapped down to $\ell_{t+1}$. It so happens that this non-monotonicity—even at a single iteration—can cause loss of precision. That is, problems occur if
@@ 4.5   The algorithm
We give the complete pseudocode for our proposed method for any degree in Algorithm 2. We give the specfic version of the Polar Express with degree $d=5$ and $\ell=10^{-3}$ used in our GPT experiments in Algorithm 1. Our algorithm first computes the polynomials $p_{1},\ldots,p_{T}$ of Theorem 4.1 in full precision using the results of Section 4.2 (or the Remez algorithm for $d>5$). This stage is offline because the coefficients of the polynomials are only computed and stored once. For every subsequent call to the algorithm, these coefficients are reused and the offline stage is skipped. For instance, in Algorithm 1 these polynomials have been precomputed and stored in the variable coeffs_list.
@@
output: An approximation $\bm{X}_{T}$ to $\operatorname*{polar}(\bm{M})$.
1:Offline stage: $\triangleright$ In float64
2:$\ell_{1}=\ell$, $u_{1}=1$.
3:for $t=1,2,\ldots,T$ do
4:     Solve using Remez Appendix D:
5:     $p_{t}=\operatorname*{argmin}\limits_{p\in\mathbb{P}_{d}^{\operatorname*{odd}}}\max\limits_{x\in\left[\max(\ell_{t},u_{t}/10),\,u_{t}\right]}|1-p(x)|$.
6:     $p_{t}\leftarrow p_{t}(\cdot/1.01)$.
7:     $\ell_{t+1}=p_{t}(\ell_{t}),u_{t+1}=2-\ell_{t+1}$.
8:end for
9:Online stage: $\triangleright$ In float16
10:Let $\bm{X}_{0}=\bm{M}/(\|\bm{M}\|_{\text{F}}+10^{-7})$.
11:for $t=1,2,\ldots,T$ do
12:     $\bm{X}_{t}=p_{t}(\bm{X}_{t-1})$.
13:end for
14:return $\bm{X}_{T}$.
@@ 5.1   Convergence of Polar Express
We first study an idealized scenario where the spectrum of the input matrix is known exactly. We generate a random matrix whose singular values are evenly spaced on a logarithmic scale between $10^{-6}$ and $1$. The right and left singular vectors are chosen at random. The left panel of Figure 4 shows the results. Since all the methods in this plot use degree-5 polynomials, their computational and runtime costs are all proportional to the number of iterations. As expected, Newton-Schulz converges but makes almost no progress for the first 17 iterations. Jordan’s method rapidly achieves an error of $\approx 0.3$ after just 11 iterations, but ceases to converge further. You’s method, which is difficult to see on the plot because it is only defined for six iterations, converges at a similar rate as Jordan’s method. When Polar Express is instantiated with $\ell=\sigma_{\min}$, it dominates the other methods at every iteration, achieving excellent accuracy after just 11 iterations and converging about twice as fast as Newton-Schulz to any given error. Even when the lower bound on $\sigma_{\min}$ is wrong by two orders of magnitude in either direction, the method remains competitive, though it does not actually outperform Keller until iteration 13 or 14.
Figure 5: Convergence of polynomial methods in the Frobenius norm on GPT-2 gradient matrices. The number of matrix-matrix products is $T(d+1)/2$, where $d$ is the degree ($3$ for Chen & Chow; $5$ for all others) and $T$ is the number of iterations.

### v3

@@ 3  Approximations by Compositions of Polynomials
Iterative methods for $\operatorname*{polar}(\bm{M})$ can be seen in this light. For instance, the degree-5 Newton-Schulz method uses the polynomial update $p_{t}(x)=\frac{15}{8}x-\frac{10}{8}x^{3}+\frac{3}{8}x^{5}$ for each $t=1,\ldots,T$. The composition $p=p_{T}\circ\cdots\circ p_{1}$ approximates $\operatorname*{sign}(x)$, and the approximation error goes to 0 as $T$ grows. In this paper, we ask the following question: what choice of $p_{T}\circ\cdots\circ p_{1}$ gives the best approximation to $\operatorname*{sign}(x)$?
@@
In other words, the problem given in (6) reduces to that of finding a “uniform” or “minimax” approximation to the constant function $x\mapsto 1$ over the interval $\ell,u]$, as given in (7). Uniform approximation on an interval by polynomials or rational functions of a given degree is a central topic in approximation theory; see e.g. [50]. Here, we seek an approximation of a particular form—a composition of odd polynomials of fixed degrees. In the next section, we solve the optimization problem of (7) and use the solution to create Polar Express. Figure 2 (a) shows the resulting $p^{*}$ polynomial labeled as PolarExp, as compared to the Jordan’s method in [25], and the six iterations of You’s method in [10].
@@ 4.1  Greedy is optimal
The key observation is that the polynomial used in each iteration can be chosen greedily, given the choice of polynomials from the previous iterations. For the first iteration, we choose $p_{1}$ so as to map the interval $\ell,u]$ as close to $1$ as possible. That is, it minimizes $\max_{x\in[\ell,u]}|1-p_{1}(x)|$. The image of $p_{1}$ will be a new interval $\ell_{2},u_{2}]$, where
@@
We now pick $p_{2}$ to map the interval $\ell_{2},u_{2}]$ as close to $1$ as possible, obtaining a new interval $\ell_{3},u_{3}]$ that is the image of $\ell,u]$ through $p_{2}\circ p_{1}$. We continue this process for as many iterations as desired.
@@
The resulting composition $p^{\star}:=p_{T}\circ p_{T-1}\circ\cdots\circ p_{1}$ is optimal and the error is given by:
Furthermore the new error, lower and upper bounds can be computed through
@@
Our theory translates entirely to this more general case. However, for simplicity we assume $d=d_{t}$ for all $t=1,\ldots,T$. Our setting is similar to that of [32], which considers the closely related problem of choosing the depth $T$ and degrees $d_{1},\ldots,d_{T}$ such that $p$ approximates $\operatorname*{sign}$ up to a prescribed error tolerance while minimizing the number of scalar multiplications. Interestingly, from [32, Table 2] the optimal choice of degrees is $d_{t}=5$ for almost all iterations. This justifies choosing $d$ to be a constant and our use of $d=5$ in particular.
Fortunately, (11) shows that once $p_{t}$ has been found, we can compute the new lower and upper bounds $\ell_{t+1}$ and $u_{t+1}$ and the approximation error simply by evaluating $p_{t}(\ell_{t})$. Hence, for any fixed upper and lower bounds on the singular values of $\bm{M}$, we can precompute the polynomials $p_{1},\ldots,p_{T}$ and the bounds $\ell_{1},u_{1}],\ldots,[\ell_{T+1},u_{T+1}]$. Then, applying the iterative procedure of (5), the final iterate $\bm{X}_{T}$ will satisfy the following error bound
@@
In fact, Theorem 4.3 underestimates how fast our method converges. For degree $d=5$, our method converges about twice as fast as Newton-Schulz (compare with [12, Section 3.1]). Furthermore, the same analysis applies even if $p^{*}$ is constructed using a “lower bound” $\ell$ that was too high. That is, replacing $\ell$ on the right-hand side of (13) by $\sigma_{\min}$, the theorem holds even if $p^{*}$ is constructed to be optimal on the interval $\ell,1]$ for $\ell>\sigma_{\min}$. Intuitively, when $\ell=u=1$, the polynomial $p^{*}$ coincides exactly with the Newton-Schulz method. Mistakenly setting $\ell>\sigma_{\min}$, we obtain a polynomial that converges slower than the optimal polynomial but faster than Newton-Schulz, so the guarantee of Theorem 4.3 still holds (cf. [12, Theorem 3.3]).
@@ 4.2  Finding the optimal polynomial for each iteration
Theorem 4.1 shows that we can solve (7) by greedily choosing the optimal approximation $p_{t}\in\mathbb{P}_{d}^{\operatorname*{odd}}$ for each interval $\ell_{t},u_{t}]$ for $t=1,\ldots,T$. In this section, we show how to find each $p_{t}$. Since we are now focused on just one iteration, we drop the subscripts. Given $\ell$ and $u$, we wish to solve the following optimization problem:
@@
That is, we seek a minimax or uniform approximation of the function $x\mapsto 1$ on $\ell,u]$ from the set of odd polynomials. (Equivalently, we seek a minimax optimal approximation to $\operatorname*{sign}(x)$ on $u,-\ell]\cup[\ell,u]$.)
Problems of the form (14) are well-studied in approximation theory and numerical analysis. The key mathematical insight underlying the solution is the Equioscillation Theorem, which we state formally for our setting in Lemma A.1. This theorem gives a surprising characterization of the optimal solution of (14): an odd $p$ is optimal for degree $2q+1$ if and only if there is a set of $q+2$ equioscillating points. This is a set of points at which $p$ achieves its maximum approximation error $\pm E$, and for which the sign of the error alternates. Even if the optimal approximation error $E$ is not known in advance, finding a set of $q+2$ equioscillating points for a given $E$ serves as a certificate that no better approximation error is achievable. The Equioscillation Theorem is the basis of the Remez algorithm [42, 43], a general tool that can be used to find (nearly) optimal polynomial approximations of a given degree to any function on any interval. With very minor modifications to handle the constraint that $p$ be odd, Remez can be used to directly solve (14).
@@
We briefly describe the solution for $d=3$. We seek a polynomial of the form $p(x)=ax+bx^{3}$. The Equioscillation Theorem stipulates that $p$ must have an equioscillating set of size 3. For $p$ to achieve its maximum error at a point $x$, $x$ must be a local extremum of p(x)1p(x)-1 on the interval $\ell,u]$. Thus, for $x$ to be eligible for membership in the equioscillating set, it must either be a true local extremum of p(x)1p(x)-1 that happens to lie in $\ell,u]$, or else one of the endpoints $\ell,u$. However, because $p$ is an odd cubic, it has at most one true local extremum on $\mathbb{R}_{\geq 0}$. Thus, to build an equioscillating set of three points, we must include $p$’s unique positive local extremum and both endpoints. This local extremum of $p$ occurs at $\sqrt{\frac{-a}{3b}}$. Therefore, we seek $a,b$ such that
@@
We now turn to the degree-5 case. The intuition of Algorithm 3 is as follows. For any fixed set of four points $\ell<q<r<u$, we can find a degree-5 odd polynomial $p(x)=ax+bx^{3}+cx^{5}$ that satisfies
@@
for some $E$ by solving a $4\times 4$ linear system in $a,b,c$ and $E$. Likewise, for any fixed degree-5 odd $p$, we can find its four (or fewer) local extrema on $\ell,u]$ as follows: they occur at $\ell,u$ and the roots of $p^{\prime}$, which is an even degree-4 polynomial whose roots can easily be found by the quadratic formula. Algorithm 3 simply alternates between these two steps (solving for $a,b,c,E$ and solving for $q,r$) until the points $q,r$ converge. Once they have converged, $\{\ell,q,r,u\}$ forms an equioscillating set, so $p$ is the optimal polynomial. For more details, please see Appendix D.
@@ 4.4  Finite precision considerations
To fix this issue, we simply replace each polynomial $x\mapsto p_{t}(x)$ by $x\mapsto p_{t}(x/1.01)$. This safety factor corrects for round-off errors in previous iterations while only slightly changing the behavior of the polynomial on the interval $\ell_{t},u_{t}]$, though it does cause the singular values to converge to $0.999998$ instead of to $1$. To correct for this, the safety factor can be omitted in the final iteration.
The second issue was identified in [39] and addressed in the context of polynomial iterations by Chen and Chow [12]. In general, iterative methods for $\operatorname*{polar}(\bm{M})$ aim to increase each singular value relative to the largest singular value; while $\sigma_{\min}(\bm{X}_{0})\ll\sigma_{\max}(\bm{X}_{0})$, after enough iterations, $\sigma_{\min}(\bm{X}_{t})\approx\sigma_{\max}(\bm{X}_{t})\approx 1$. However, the convergence of each singular value to $\sigma_{\max}$ may not be monotonic. Over the domain $\ell_{t},u_{t}]$, our optimal polynomial $p_{t}$ oscillates repeatedly between $\ell_{t+1}$ and $u_{t+1}$, so some singular values that are near $u_{t}$ may get mapped down to $\ell_{t+1}$. It so happens that this non-monotonicity—even at a single iteration—can cause loss of precision. That is, problems occur if
@@ 4.5  The algorithm
output: An approximation $\bm{X}_{T}$ to $\operatorname{polar}(\bm{M})$.
1
2 $\ell_{1}=\ell$, $u_{1}=1$.
3for $t=1,2,\ldots,T$ do
4  Solve using Remez (Appendix D):
5  $p_{t}=\operatorname*{argmin}\limits_{p\in\mathbb{P}_{d}^{\operatorname*{odd}}}\max\limits_{x\in\left[\max(\ell_{t},u_{t}/10),\,u_{t}\right]}|1-p(x)|$
6  $p_{t}\leftarrow p_{t}(\cdot/1.01)$
7  $\ell_{t+1}\leftarrow p_{t}(\ell_{t})$, $u_{t+1}\leftarrow 2-\ell_{t+1}$
8end for
9
10
11 Set $\bm{X}_{0}=\bm{M}/(\|\bm{M}\|_{\text{F}}+10^{-2})$.
12for $t=1,2,\ldots,T$ do
13  $\bm{X}_{t}=p_{t}(\bm{X}_{t-1})$
14end for
15return $\bm{X}_{T}$.
Offline: precompute polynomials in float64 Online: apply cached polynomials in bfloat16
We give the pseudocode of our proposed method for any degree in Algorithm 2 and provide a complete implementation in our repository. We give the specific version of the Polar Express with degree $d=5$ and $\ell=10^{-3}$ used in our GPT experiments in Algorithm 1. Our algorithm first computes the polynomials $p_{1},\ldots,p_{T}$ of Theorem 4.1 in full precision using the results of Section 4.2 (or the Remez algorithm for $d>5$). This stage is offline because the coefficients of the polynomials are only computed and stored once. For every subsequent call to the algorithm, these coefficients are reused and the offline stage is skipped. For instance, in Algorithm 1 these polynomials have been precomputed and stored in the variable coeffs_list.
@@ 5.1  Convergence of Polar Express
We first study an idealized scenario where the spectrum of the input matrix is known exactly. We generate a random matrix whose singular values are evenly spaced on a logarithmic scale between $10^{-6}$ and $1$. The right and left singular vectors are chosen at random. The left panel of Figure 4 shows the results. Since all the methods in this plot use degree-5 polynomials, their computational and runtime costs are all proportional to the number of iterations. As expected, Newton-Schulz converges but makes almost no progress for the first 17 iterations. Jordan’s method rapidly achieves an error of $\approx 0.3$ after just 11 iterations, but ceases to converge further. You’s method, which is difficult to see on the plot because it is only defined for six iterations, converges at a similar rate as Jordan’s method. When Polar Express is instantiated with $\ell=\sigma_{\min}$, it dominates the other methods at every iteration, achieving excellent accuracy after just 11 iterations and converging about twice as fast as Newton-Schulz to any given error. Even when the lower bound on $\sigma_{\min}$ is wrong by two orders of magnitude in either direction, the method remains competitive, though it does not actually outperform Jordan’s method until iteration 13 or 14.
Figure 5: Convergence of polynomial methods in the Frobenius norm on GPT-2 gradient matrices. The number of matrix-matrix products is $T(d+1)/2$, where $d$ is the degree (33 for Chen & Chow; 55 for all others) and $T$ is the number of iterations.
@@
Figure 6: Training a GPT-2 (124M) model on 1 Billion tokens of the Fineweb data set [2]. The Legend muon-<name> refers to using muon with the <name> method for computing $\operatorname*{polar}(\bm{M})$ with weight decay zero. Top Left: The final validation loss vs. the learning rate. The final best validation losses for each method were, in reverse order, adamw: $4.197$, muon-Jordan: $3.639$, muon-You: $3.629$ and muon-PolarExp: $3.588$. Bottom Left: The final training loss vs the learning rate. Top Right: Validation loss vs. number of iterations. Bottom Left: validation loss vs. time, plotting each method with its best learning rate.
Figure 7: Training a GPT-2-Large model (774M params) on 1 Billion tokens of the Fineweb data set [2]. The Legend muon-<name> refers to using muon with the <name> method for computing $\operatorname*{polar}(\bm{M})$. We used weight decay $0.1$ for all methods. Top Left: The final validation loss vs. the learning rate. The final best validation losses for each method were, in reverse order, adamw: $4.197$, muon-Jordan: $3.639$, muon-You: $3.629$ and muon-PolarExp: $3.588$. Bottom Left: The final training loss vs the learning rate. Top Right: Validation loss vs. number of iterations. Bottom Left: validation loss vs. time, plotting each method with its best learning rate.